OpenStudy (anonymous):

LINEAR ALGEBRA. verifying (k+m)(u)=ku+mu... where k,m are any integers and u is an element of V={(x,y,z) | z=3}. addition defined by (x1,y1,z1)+(x2,y2,z2) = (x1+x2,y1+y2,3) scalar multiplication defined by s(x,y,z)=(sx,sy,3)

OpenStudy (anonymous):

I am learning linear algebra as well, can you tell me what topic this comes from ?

OpenStudy (anonymous):

determining is v is a vector space... so we have to verify the ten axioms

OpenStudy (anonymous):

so we are finding what order space this is?

OpenStudy (anonymous):

mm.. what u mean with order space?

OpenStudy (anonymous):

line plane is 2nd order space, line is 1st order space

OpenStudy (anonymous):

ohhh,,, well my example is already i 3 space

OpenStudy (anonymous):

if u already covered this section.. we defined vector space as a set of vectors on which are defined by two operations: 1) vector addition 2) vector multiplication and must satisfy 10 axioms

OpenStudy (anonymous):

2)scalar multiplication***

OpenStudy (anonymous):

I know about first two , but never heard of this '10 axioms' stuff

OpenStudy (anonymous):

are you learning it independently or with class?

OpenStudy (anonymous):

class

OpenStudy (anonymous):

1. u + v 2 V (V is closed under addition) 2. u + v = v + u (commutative property) 3. (u + v) + w = u + (v + w) (associative property) 4. There exists a vector in V , called the zero vector and denoted 0 such that u + 0 = u (additive identity) 5. For every vector u in V , there exists a vector 􀀀u also in V such that u + (􀀀u) = 0 (additive inverse) 6. cu 2 V (V is closed under scalar multiplication) 7. c (u + v) = cu + cv 8. (c + d) u = cu + du 9. c (du) = (cd) u 10. 1u = u

OpenStudy (anonymous):

so in this case im trying to verify number 8

OpenStudy (anonymous):

distributive property

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

do you watch MIT lectures

OpenStudy (anonymous):

mmm nop

OpenStudy (anonymous):

are u learning independenly

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

I tend to do better in class , if I already learned it beforehand

OpenStudy (anonymous):

have you done differential equation yet?

OpenStudy (anonymous):

yes in calculus..

OpenStudy (anonymous):

im learning it now

OpenStudy (anonymous):

let me know if you need any help with differential equation; I am really good at it

OpenStudy (anonymous):

:).. cool... i like derivative for some reasons... where im having a little trouble is in double and triple integrals... are u also learning that independnly

OpenStudy (anonymous):

no, I just finished the class this semseter

OpenStudy (anonymous):

so u r good with double and triples integrals too?

OpenStudy (mathmate):

Given u(x,y,z), (x1,y1,z1)+(x2,y2,z2) = (x1+x2,y1+y2,3) s(x,y,z)=(sx,sy,3) Assuming vector space is defined over field F such as R. (k+m)(u) =(k+m)(x,y,3) [definition of u] =((k+m)x,(k+m)y,3) [definition of scalar multiplication] =(kx+mx, ky+my,3) [distributivity defined over F] =(kx,ky,3)+(mx,my,3) [definition of addition] =k(x,y,3)+m(x,y,3) [definition of scalar multiplication] =ku+mu [definition of u] QED